OpenStudy (anonymous):

I really don't get this 1.) f(n)=4n+1 g(n)=4n find ( f o g )(6) ^f of g its not times 2.) g(x)=3x-4 h(x)=x+2 Find ( g . h)(-8) 3.) g(x)=x^3 - 2x h(x)=3x+2 Find (2g-4h)(2)

OpenStudy (anonymous):

Are you saying (f(g))?

OpenStudy (bahrom7893):

yes that's what it means f(g(6))

OpenStudy (bahrom7893):

1) f(n)=4n+1 g(n)=4n find ( f o g )(6) g(6) = 4*6=24 f(g(6)) = 4*24+1 = 97 etc

OpenStudy (anonymous):

If so, then #1 would be 4(4(6)+1 = 97 #2 3(-6) -4 = -24

OpenStudy (anonymous):

Oh look i'm blue now :D #3 4x^3 - 40x - 16

OpenStudy (bahrom7893):

im not sure about 2 and 3

OpenStudy (anonymous):

Omg you guys thank you

OpenStudy (bahrom7893):

wait

OpenStudy (bahrom7893):

2.) g(x)=3x-4 h(x)=x+2 Find ( g . h)(-8) is that also g(h(x))?